OpenStudy (anonymous):

wy is an altitude of XYZ. What is the length of wx? the length of the altitude (wy) is 12. the length of wz is 22.

OpenStudy (amistre64):

12 tall; and 22 for the other side of the base..... .... we can find a length for yz; but is there any other info for this?

OpenStudy (amistre64):

yz = sqrt(12^2 + 22^2) and that can give us an angle too maybe

OpenStudy (amistre64):

tan(z) = 12/22 z = tan-1(12/22)